Systems Staff Meeting Notes 3/20/00

Systems Staff Meeting Notes
March 20, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. An order is on the way - a tiny package with what looks like 6 grains of sand in a big box. In actuality, they are bearings from Logitech.
  2. All parts for the Apache server (except the motherboards) have been received. We'll need 5 1/4 mounting kits - Everett & Lindsay will race to see who can assemble it the fastest.
Liaisons
  1. Liaison announcement - Con Con crash - command parser bug. Microsoft has a patch out now - it is currently available on NT2. It seems to have no other effects. Not for use on NT machines.
  2. Virus alert: W32 Melting virus - it is a worm-type virus. It can modify your HD and eat things. It's not up on McAfee yet. They have had a dat file upgrade, hopefully the Melting virus scan will be up this week.
Networking
  1. NT10 will replace the Novell server. We are currently installing a demo of Juke Box software on it.
  2. Wireless - either the wiring or more antennas are needed. A replacement PCMCIA for the laptop as well as a new battery are also required.
  3. Novell to NT - CD shortcuts, etc. will have to be changed.
  4. IPX access to NT - as soon as the hardware is ready, we'll test. There will be a small amount of down time - a major changeover is required, and there are some festering problems with the version of Novell that we have.
  5. Links to CD resources from Web page - in general, they are running specific applications for access. We'll need to look at it.
  6. OS running on the Web - metaOS level will supposedly run on all servers. Will check into multi-platform based on web extensions. They are fairly costly & another drawback is that you have to use its internal database.
  7. Switched port changeover for staff - need to know how many switches are necessary to replace all hubs. An estimated is needed - 1900's will be traded in. Space issue - recycled bathroom, Grace's old room - we'll check on the viability of moving the servers.
  8. Cable tray installation in West - stopped for lack of permits - should resume on Tuesday. Progress will be checked by Will.
Software Projects Status
  1. Q&A database - demo scheduled for the end of the month. It's close to completion.
  2. Library Statistics - prototype will be demoed today.
  3. Newspaper project - data has been migrated to SQL - first cut this week. The URL has been given to the outside world.
  4. Systems wiring database - will document at least one room this week.
  5. LASSIE - Hilton will start on it in April.
  6. Online equipment reserve - need to discuss when Debra is present.
Calendar
  1. Ying out March 24.
  2. Will out April 3-7.
  3. Debra out April 18-21.


Back to the Systems Homepage

Last updated March 27, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu